OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2).

OpenStudy (paxpolaris):

slope intercept form:\[\Large y=mx+b\] where m is slope, & b is y-intercept....{i.e. b is y-value ...when x=0(graph hits y-axis)}

OpenStudy (paxpolaris):

so, you can get b by plugging in (x,y)=(0,-2)\[-2=\cancel{m \times0}+b\]\[\implies \Large b=-2\]

OpenStudy (paxpolaris):

\[m=slope={rise \over run}=rate\ of\ change = { change\ in\ y \over change\ in\ x}\]

OpenStudy (paxpolaris):

\[m={-2-8 \over 0-4} = {10 \over -4} =\Large \frac 52\] \[\Large \therefore\ y= \frac 52 x -2\]